Dickson quantity demand of perfume rice is increased from 34bags to 67 bags when his income increased from $100 to $500.calculate the income elasticity of demand. Is the perfume rice normal or inferior? Explain
1
Expert's answer
2020-04-26T18:58:38-0400

D1=34

D2=67

I1=$100

I2=$500

"\text{income elasticity of demand}=\frac{(D2-D1)/(D2+D1)}{(I2-I1)/(I2+I1)}"

=(33/101)/(400/600)= 0.327/0.667=0.49